OpenStudy (anonymous):

Find the equation of the circle that has: endpoints of the diameter: (2,5) and (4,-1)

hero (hero):

Hint: Find the midpoint of the diameter.

OpenStudy (anonymous):

wat was the midpoint equation again? xD haha..

hero (hero):

Really bro? I mean, really? Do I have to spell everything out for you every time? Some things you should be able to figure out on your own.

OpenStudy (anonymous):

hahaok ill check on wikipedia

hero (hero):

\[\left(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\right)\]

OpenStudy (anonymous):

ok for the midpoint i got (-1,2)

hero (hero):

That is not correct.

OpenStudy (anonymous):

oh wait sorry its late haha i meant (3,2) :)))))))

hero (hero):

So that's the center of your circle. Now find the radius.

OpenStudy (anonymous):

so i use distance formula using these points?

OpenStudy (anonymous):

my teacher got (x-3)^2 + (y-2)^2 = 10 shouldn't it be 25?

OpenStudy (anonymous):

the previous question of this said that the radius is 5

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@binary3i can u help me? ^^

OpenStudy (binary3i):

you know the end points of diameter so mid point of diameter will be the center of circle can you do that? and radius is the half of diameter so can you also find that using distance between the end points of diameter then dividing it by 2. then using the general equation

hero (hero):

@lilfayfay, if the answer was r^2 = 10, then you can assume that the radius was sqrt(10)

hero (hero):

Then further prove it to yourself by actually doing the work.

OpenStudy (anonymous):

well i get how my teacher got (x-3)^2 + (y-2)^2 but it dont get the (x-3)^2 + (y-2)^2 = 10

OpenStudy (anonymous):

@binary3i so how did he get =10??

hero (hero):

I just told you how. :/

OpenStudy (anonymous):

no it was r^2=25 from the previous question center at (2, -4) and radius of 5.

OpenStudy (anonymous):

or does the previous question have nothing to do with this question

OpenStudy (binary3i):

(x-h)^2 + (y-k)^2=R^2 where h and k are the midpoints and R is the radius R is from dividig the diameter by 2 did you used the distance formula for the end points of the diameter

OpenStudy (anonymous):

so i do (x-3)^2 + (y-2)^2 = R^2?

hero (hero):

Yes bro. Geez, do you pay attention to any of these formulas?

OpenStudy (anonymous):

@binary3i so what do i do next

hero (hero):

Bro, I explained it to you. When the other guy explains it, you will see that it is r = sqrt(10) r^2 = 10

OpenStudy (anonymous):

but myteacher put (x-3)^2 + (y-2) = 10 and wheredid the 10 come from? From the previous question

OpenStudy (binary3i):

\[2R=\sqrt{(4-2)^2 + (-1-5)^2}=\sqrt{2^2 + 6^2)}=\sqrt{4+36}=\sqrt{40}=2\sqrt{10}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

im just finding the equation of the circle

hero (hero):

Divide that by 2 to get \sqrt(10)

OpenStudy (anonymous):

so the equation of the circle is:\[(x-3)^2 + (y-2)^2 = \sqrt{10}\]?

hero (hero):

No bro For the 10000th time \[r = \sqrt{10}\] \[r^2 = 10\] The formula needs to be written in terms of \(r^2\). Get it?

OpenStudy (anonymous):

but the question is only asking me to write the equation of the circle not find the radius BRO

hero (hero):

Bro, in order to write the equation of the circle, you need the square of the radius.

OpenStudy (anonymous):

yeah so its (x-3)^2 +(y-2)^2 =sqr10 BROOooooooooooo

hero (hero):

Bro, it's 10, not sqrt(10). Now you're just being mental.

OpenStudy (anonymous):

how is it 10........

hero (hero):

Now you want me to explain it to you 10,001 times

OpenStudy (binary3i):

see R = sqrt{10} R^2=10 eqution of circle is (X-h)^2 + (Y-k)^2= R^2

OpenStudy (anonymous):

lol :P xD

hero (hero):

@lilfayfay, @binary3i agrees with me.

hero (hero):

He just wrote the exact same thing I wrote \[r = \sqrt{10}\] \[r^2 = 10\]

OpenStudy (binary3i):

yes hero i stand with you

OpenStudy (anonymous):

so the answer is \[\sqrt{(x-3)^2 + (y-2)^2} = 10\]

hero (hero):

Yes, finally, after all these years....you figure it out. Even though all you needed was two seconds.

OpenStudy (anonymous):

why didnt u tell me earlier that I left out the square root

hero (hero):

Good question.

OpenStudy (anonymous):

............................ :P

hero (hero):

Hey bro, I can't correct every single mistake you make. Some things you have to figure out on your own.

OpenStudy (anonymous):

no wait it's just (x-3)^2 + (y-2)^2 = 10

hero (hero):

Confusing yourself again.

OpenStudy (anonymous):

no that's correct, thats what my teacher got

hero (hero):

Your teacher doesn't make mistakes?

OpenStudy (anonymous):

he does but thats the answer he gave and reviewed with the whole class

hero (hero):

Okay bro. Good luck with that. By the way, when I was in school, I corrected my teacher on numerous occasions. If I came to your school, I would correct your teacher as well. Fortunately, I'm not a student anymore. But you are, and when you have the opportunity to correct your teacher, you should. All the students in your class must be afraid of your teacher.

OpenStudy (anonymous):

nope it's completely the opposite ^^

OpenStudy (anonymous):

lol